If compound interest for two successive years is Rs.110 and Rs.121 respectively, then the rate of interest is

Grade:12th pass

1 Answers

Vikas TU
4 years ago
Rs 110 and 121 is the compound interest 
Additional interest = 121-110 = 11 in the second year.
So,  Rate of interest = 11*100/(110*1)
This will be equal to 10%